OpenStudy (anonymous):

If two cards are chosen at random from a standard deck of playing cards, what is the probability that at least one is an Ace or a King? Round your answer to two decimal places.

OpenStudy (amistre64):

2/13 perhaps?

OpenStudy (amistre64):

ugh ... at least ..

OpenStudy (anonymous):

idk?

OpenStudy (amistre64):

itd help if you had some idea :)

OpenStudy (amistre64):

at least one ... means at most 2 .

OpenStudy (amistre64):

maybe its better to see it as at most none ...

OpenStudy (anonymous):

yes that is the trick. compute the probability that neither are. then subtract from 1

OpenStudy (amistre64):

52 -8 --- 44 1 - 44/52 + 43/51 then?

OpenStudy (anonymous):

neither are is easy. first one isn't has probability \[\frac{44}{52}\] then second one isn't given that the first one isn't has probability \[\frac{43}{51}\] hope the reasoning is clear

OpenStudy (anonymous):

Well there's 52C2 ways to pick 2 cards, and 44C2 ways to not pick an A or K. So your solution would be 1-(44C2/52C2) or something.

OpenStudy (amistre64):

typoed meself lol

OpenStudy (anonymous):

so probability neither are is \[\frac{44}{52}\times \frac{43}{51}\] and your answer is \[1-\frac{44}{52}\times \frac{43}{51}\]

OpenStudy (anonymous):

i am sure there is another way to do it, but it will come out to this answer.

OpenStudy (anonymous):

oh no never mind. forget that

OpenStudy (amistre64):

does the "or" mean addition? or is that beside the point?

OpenStudy (amistre64):

P(-A or -B)

OpenStudy (anonymous):

.29?

OpenStudy (anonymous):

that is what i got if you round

OpenStudy (anonymous):

@mooby your answer is the same as mine. cancel the 2 in top and bottom and you get \[1-\frac{44\times 43}{52\times 51}\]